# What’s Wrong with Your Code Generated by Large Language Models? An Extensive Study

### v1

@@ 8. Threats to Validity
Dataset Validity. The first concern involves potential issues with the popular benchmarks we used. APPS+ contains a wide range of programming problems with varying difficulty, which can thoroughly test the code generation capabilities of LLMs. However, there are certain issues with this benchmark. First, the problem descriptions often include a lot of background information irrelevant to the core algorithm, which can interfere with the LLM’s understanding of the problem. Furthermore, some descriptions use images for explanations, but these images are merely represented by “[image]”, which prevents the LLM from fully comprehending the problem. To avoid these issues, we manually checked APPS+ and carefully removed cases with inaccurate or incomplete descriptions. Another issue is that all benchmarks we used were monolingual, i.e., the programming language of canonical code was in Python. In practice, current commonly used benchmarks including HumanEval+ (Chen et al., 2021), MBPP+ (Austin et al., 2021) and APPS+ (Hendrycks et al., 2021) (these benchmarks also being those which we used in our study) exclusively feature Python, and researchers also focus on enhancing LLM’s code generation capabilities in Python (Contributors, 2023; Bai et al., 2023; Liu et al., 2024b; Zhuo et al., 2024). Consequently, the results of LLMs on these benchmarks are indicative of their overall code generation proficiency, and our findings can be generalized for application in other experimental settings. In the future, we plan to investigate and analyze limitations faced by LLMs in generating other programming languages. The final issue is that the benchmarks may cause data leakage (Jain et al., 2024; Golchin and Surdeanu, 2023; Weller et al., 2023; Riddell et al., 2024; Roberts et al., 2023; Shi et al., 2023; Zhou et al., 2023) as current LLMs are trained on a large fraction of GitHub, StackOverflow and code libraries (Chen et al., 2021). In our process of benchmark construction, we collected samples from the latest code repositories and utilized a de-duplication algorithm (i.e., MinHash (Broder, 1997)) on code training corpora including open-source bases (i.e., RedPajama Data v2, Stack v1 (Kocetkov et al., [n. d.]) and v2 (Lozhkov et al., 2024), RefinedWeb (Penedo et al., 2023), and Pile (Gao et al., 2020)) and our internal code training corpora, to mitigate data leakage.
External Validity. Threats to external validity relate to the ability of our results to be generalized according to different scenarios. We address these concerns by selecting a diverse set of seven state-of-the-art LLMs, including two open-source textural LLMs, two open-source code LLMs, and three closed-source LLMs. These LLMs are widely used in a variety of fields. Regarding benchmarks, we use three well-known datasets. We also carefully construct a real-world benchmark, RWPB, to evaluate the performance of LLMs on real-world code generation. Our experimental setting of LLMs adheres to protocols established in prior works, which ensures that our results and findings can be generalized for application in other scenarios.
@@
Internal Validity. When labeling bug types in code generation, we use manual categorization with the assistance of GPT-4. This approach can improve the accuracy of the process but may cause “lazy annotation” (details in Section 7). Additionally, different experts may have varying understandings of bug types, which potentially might have resulted in labeling errors. To ensure consistency, we employ double-checking among different experts. This method helps us achieve a consensus on bug types, avoiding incorrect labeling. Furthermore, when faced with the same problem, LLMs may generate differing responses. To ensure the consistency of responses, we set the temperature to 0.1 and top-k to one, in accordance with prior common settings (gpt, 2023b; Guo et al., 2024).
9. Related Work
@@
Large Language Models for Code Generation. Recently, the advancement of LLMs has significantly propelled the field of code generation. Due to training on large code corpora, several text LLMs have demonstrated remarkable ability in understanding natural language and synthesizing code, as seen in models such as Mistral (Jiang et al., 2023), Qwen1.5 (Bai et al., 2023), DeepSeek-LLM (DeepSeek-AI, 2024), the ChatGPT family (gpt, 2023a; OpenAI, 2023), the Claude3 family (Anthropic, 2024), and the LLama family (Touvron et al., 2023; lla, 2023, 2024). On the other hand, code LLMs, primarily trained on extensive code data including CodeGPT (Lu et al., 2021), SantaCoder (Allal et al., 2023), CodeGeex (Zheng et al., 2023b), CodeX (Chen et al., 2021), WizardCoder (Luo et al., 2023), CodeLlama (Roziere et al., 2023), DeepSeekCoder (Guo et al., 2024), Phi-3-Instruct (Abdin et al., 2024), and StarCoder-2-Instruct (Li et al., 2023), have also exhibited remarkable capabilities in code generation. Researchers have further improved the performance of LLMs on code generation by prompt engineering (Le et al., 2023; Shin et al., 2023; Denny et al., 2023) or by introducing supervised fine-tuning (Jain et al., 2023; Bairi et al., 2023; Weyssow et al., 2023) and reinforcement learning (Dou et al., 2024; Yu et al., 2023; Le et al., 2022; Liu et al., 2023b; Shojaee et al., 2023; Zheng et al., 2023a).
Bug Study in Software Engineering. The investigation of bugs helps developers comprehensively understand the code tasks and the defects of models, and introduce improvement approaches based on findings. Several works analyze bugs introduced in deep learning models and frameworks (Islam et al., 2019; Wang et al., 2022b; Zhang et al., 2018; Garcia et al., 2020). Additionally, techniques based on LLMs have demonstrated impressive abilities in coding tasks. Thus several works have analyzed bugs introduced by LLMs in a variety of coding tasks (Pan et al., 2024; Xia and Zhang, 2022; Yang et al., 2024; Liu et al., 2024a; Mousavi et al., 2024; Tambon et al., 2024). In comparison to these studies, we studied bugs introduced by LLMs in code generation. Tambon et al. (Tambon et al., 2024) analyzed bugs introduced by LLMs while generating non-standalone code, which differs from our study (Tambon et al., 2024). Moreover, the LLMs used in this study are outdated, which may lead to experimental findings that do not fit into current commonly used LLMs. They also lack an analysis of data leakage when constructing the benchmark, as the data may have been seen in the training phase of LLMs, which would lead to inaccurate results. In contrast to this study, we evaluated the code generated by LLMs from multiple perspectives, understood bugs introduced by these models and introduced some direct methods to improve the code generation ability of LLMs based on our findings. We also elaborately constructed a real-world benchmark while working to avoid data leakage, and comprehensively investigated the capability of these models on code generation for real-world projects.
@@
Code Generation Benchmarks. Several works have proposed benchmarks to evaluate the capability of LLMs in code generation. Some benchmarks are constructed from code exercises and introductory programming problems (Chen et al., 2021; Austin et al., 2021; Liu et al., 2024b), or completion questions (Hendrycks et al., 2021; Dai et al., 2024). However, these benchmarks are far removed from real-world program development. Some benchmarks are collected from the popular public libraries and code repositories (Zan et al., [n. d.]; Wang et al., 2022c), which may cause data leakage due to significant overlap with the datasets used for training LLMs (Li et al., 2023; Chen et al., 2021). Furthermore, several researchers aim to introduce non-standalone (Iyer et al., 2018; Yu et al., 2024), repository-level (Liu et al., 2023a; Li et al., 2024), and GitHub issue (Jimenez et al., 2023) benchmarks. In contrast to these benchmarks, our proposed RWPB aims to evaluate the code generation capability of LLMs in real-world projects and standalone scenarios. Moreover, during the filtering phase of constructing RWPB, we de-duplicated data with large open-source and our private code corpora to avoid data leakage. We also evaluated the distributions of bugs introduced by LLMs based on our constructed bug taxonomy, which further helps us understand how they perform in real-world code generation scenarios.
10. Conclusion
@@
In this paper, we conduct a comprehensive empirical investigation into the effectiveness and limitations of code generation using large language models (LLMs). We first conduct extensive experiments from multiple perspectives on seven widely-used LLMs across three widely used benchmarks. Our results reveal that these LLMs struggle to generate accurate code for those more complex problems. Subsequently, we manually annotate bug types in code generated by LLMs, construct a taxonomy of these bugs, analyze their distributions, and summarize 14 findings leading LLMs to generate erroneous code. To evaluate the effectiveness of LLMs in real-world projects, we design a rigorous benchmark construction process to minimize data leakage and construct a real-world project benchmark, RWPB. Furthermore, to mitigate bugs during code generation by LLMs, we propose a novel method that introduces self-critique, enabling LLMs to iteratively critique their generated codes and fix bugs.

### v2

@@ 7.5 Threats to Validity
Dataset Validity. The first concern involves potential issues with the popular benchmarks we used. APPS+ contains a wide range of programming problems with varying difficulty, which can thoroughly test the code generation capabilities of LLMs. However, there are certain issues with this benchmark. First, the problem descriptions often include a lot of background information irrelevant to the core algorithm, which can interfere with the LLM’s understanding of the problem. Furthermore, some descriptions use images for explanations, but these images are merely represented by “[image]”, which prevents the LLM from fully comprehending the problem. To avoid these issues, we manually checked APPS+ and carefully removed cases with inaccurate or incomplete descriptions.
Another issue is that all benchmarks we used were monolingual, i.e., the programming language of the canonical code was in Python. Our study focuses exclusively on Python rather than conducting experiments across multiple programming languages for several reasons: (1) current popular LLMs demonstrate their strongest code generation abilities in Python [74, 14]. Therefore, analyzing model performance on Python provides a reliable measure of code generation capabilities and allows for a thorough understanding of potential bugs [31]. (2) Python programming benchmarks are currently the most comprehensive, including HumanEval+ [31], MBPP+ [30] and APPS+ [15] (these benchmarks also being those which we used in our study). These benchmarks facilitate in-depth analysis of LLMs’ code generation abilities and enable detailed examination of bug types and characteristics. This is essential for understanding LLMs’ code generation performance and for broadening the scope of future research. (3) Code generation across multiple programming languages is a distinct research topic. Our study aims to provide a comprehensive understanding of LLMs’ code generation abilities and bug characteristics in Python, both in experimental and real-world settings. Including multiple programming languages would introduce considerable complexity and dilute the focus and clarity of our work. Moreover, the diversity of programming language syntax, semantics, and ecosystem-specific bugs would require a fundamentally different experimental design, which is beyond the scope of this study. Therefore, we deliberately focus on Python to ensure a rigorous and systematic investigation. In the future, we plan to investigate and analyze limitations faced by LLMs in generating other programming languages.
The final issue is that the benchmarks may cause data leakage [34, 35, 36, 37, 38, 75, 76] as current LLMs are trained on a large fraction of GitHub, StackOverflow and code libraries [31]. In our process of benchmark construction, we collected samples from the latest code repositories and utilized a de-duplication algorithm (i.e., MinHash [67]) on code training corpora including open-source bases (i.e., RedPajama Data v2, Stack v1 [64] and v2 [18], RefinedWeb [65], and Pile [66]) and our internal code training corpora, to mitigate data leakage.
External Validity. Threats to external validity pertain to the generalizability of our results across different scenarios. To address these concerns, we evaluated nine state-of-the-art LLMs, comprising four open-source textual LLMs, two open-source code LLMs, and three closed-source LLMs. We employed three established benchmarks and developed a real-world benchmark to assess LLM performance in practical code generation tasks. Our experimental setup follows protocols from prior research, supporting the generalizability of our findings to other contexts.
@@
Internal Validity. When labeling bug types in code generation, we use manual categorization with the assistance of GPT-4. This approach can improve the accuracy of the process but may cause “lazy annotation” (details in Section 7.2). Additionally, different experts may have varying understandings of bug types, which potentially might have resulted in labeling errors. To ensure consistency, we employ double-checking among different experts. This method helps us achieve a consensus on bug types, avoiding incorrect labeling. Additionally, because LLMs may generate varying responses for the same problem, we standardize the generation process by setting the temperature to 0.1 and using top-1, following established practices [39, 17].
8 Related Work
@@
Large Language Models for Code Generation. Recently, the advancement of LLMs has significantly propelled the field of code generation. Due to training on large code corpora, several text LLMs have demonstrated remarkable ability in understanding natural language and synthesizing code, as seen in models such as Mistral [77], Qwen1.5 [74], DeepSeek-LLM [78], the ChatGPT family [40, 39], the Claude3 family [41], and the Llama family [79, 44]. On the other hand, code LLMs, primarily trained on extensive code data including CodeGPT [9], SantaCoder [80], CodeGeex [81], CodeX [31], WizardCoder [82], CodeLlama [14], DeepSeekCoder [17], Phi-3-Instruct [45], and StarCoder-2-Instruct [16], have also exhibited remarkable capabilities in code generation. Researchers have further improved the performance of LLMs on code generation by prompt engineering [83, 84, 85] or by introducing supervised fine-tuning [86, 87, 88] and reinforcement learning [13, 23, 1, 62, 89, 90].
Bug Study in Software Engineering. Investigating bugs in code generation helps developers understand the limitations of models and improve their performance. Previous research has analyzed bugs introduced by deep learning models and frameworks [91, 92, 93, 94], as well as by LLMs in various coding tasks [95, 96, 97, 98, 99, 100]. In contrast to these studies, our study focuses on bugs introduced by LLMs.
We are the second study to comprehensively investigate the types of bugs that can be introduced by code generated by LLMs, following the work of Tambon et al [100]. However, Tambon et al. [100] focus on bugs in LLM-generated non-standalone code, which differs from our study. Secondly, they conducted all experiments on only one benchmark and three LLMs, which poses a threat to the validity and generalizability of their results and conclusions. In contrast, we conducted our experiments on three widely used benchmarks, one real-world benchmark, and nine popular LLMs (i.e., four open-source text LLMs, two open-source code LLMs, and three closed-source text LLMs). Thirdly, Tambon et al. categorized bugs into 10 broad categories, which we found to be quite coarse. Our categorization is more detailed, consisting of three major categories, ten secondary categories, and additional tertiary categories within each secondary category. Finally, they also overlooked the differences in bug distribution between real-world projects and existing benchmarks, as well as the potential data leakage that may occur in existing benchmarks. In contrast, we elaborately constructed a real-world benchmark while actively working to avoid data leakage. We comprehensively investigated the effectiveness of LLMs on code generation for real-world projects and analyzed LLM-generated bugs in real-world scenarios. Moreover, based on our findings, we also introduced a self-critique-based method to mitigate the bugs in LLM-generated code. In summary, our study significantly differs in terms of the types, scenarios, and sources of programming problems we focus on, as well as in our bug taxonomy and analysis perspectives. Our study encompasses a wider range of benchmarks and LLMs. Additionally, we manually constructed a benchmark of real-world scenarios, analyzed the differences in bugs within this context, and proposed a novel self-repair method based on our findings.
Code Generation Benchmarks. Several works have proposed benchmarks to evaluate the capability of LLMs in code generation. Some benchmarks are constructed from code exercises and introductory programming problems [31, 30, 49], or completion questions [15, 101]. However, these benchmarks are far removed from real-world program development. For instance, Liu et al. [49] proposed to automatically generate additional unit test samples for existing benchmarks (i.e., HumanEval and MBPP) to enhance their reliability. In contrast, we aim to construct a new high-quality real-world benchmark while making every effort to prevent data leakage to evaluate current LLMs. We also discuss the feasibility of automating our manual dataset construction process in Section 7.3.
@@
Additionally, some benchmarks are collected from the popular public libraries and code repositories [32, 22], which may cause data leakage due to significant overlap with the datasets used for training LLMs [16, 31]. Furthermore, several researchers aim to introduce non-standalone [102, 103], repository-level [104, 33], and GitHub issue [105] benchmarks. In contrast to these benchmarks, our proposed RWPB aims to evaluate the code generation capability of LLMs in real-world projects and standalone scenarios. Moreover, during the filtering phase of constructing RWPB, we de-duplicated data with large open-source and our private code corpora to avoid data leakage. We also evaluated the distributions of bugs introduced by LLMs based on our constructed bug taxonomy, which further helps us understand how they perform in real-world code generation scenarios.
9 Conclusion
@@
In this paper, we conduct a comprehensive empirical investigation into the effectiveness and limitations of code generation using LLMs. We first conduct extensive experiments from multiple perspectives on nine widely used LLMs across three popular benchmarks. Our results reveal that these LLMs struggle to generate accurate code for those more complex problems. Subsequently, we manually annotate bug types, construct a taxonomy of these bugs, analyze their distributions, and summarize 14 findings leading LLMs to generate erroneous code. To evaluate the effectiveness of LLMs in real-world projects, we design a rigorous benchmark construction process to minimize data leakage and construct a real-world project benchmark, RWPB. We analyzed the distributions of bugs on RWPB to help researchers understand how they perform in real-world code generation scenarios. Finally, to mitigate bugs during code generation by LLMs, we propose a novel method that introduces self-critique, enabling LLMs to iteratively critique their generated code and fix bugs.
[…]
The authors would like to express their sincere gratitude to Weikang ZHOU, Caishuang HUANG, Yan LIU, Enyu ZHOU, Yuhao ZHOU, Rongxiang WENG, Jingang WANG, and Xunliang CAI for their invaluable contributions to this study. The authors extend their appreciation to the annotators for their dedicated support throughout this research. The authors also wish to thank the anonymous reviewers for their insightful and constructive comments. This work was partially funded by National Natural Science Foundation of China (No.62476061, 62376061, 62206057, 62372193, U2436207), Shanghai Rising-Star Program (23QA1400200), Natural Science Foundation of Shanghai (23ZR1403500), Program of Shanghai Academic Research Leader under grant 22XD1401100.